# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ Abstract
A common obstruction to efficient sampling from high-dimensional distributions is the multimodality of the target distribution because Markov chains may get trapped far from stationarity. Still, one hopes that this is only a barrier to the mixing of Markov chains from worst-case initializations and can be overcome by choosing high-entropy initializations, e.g., a product or weakly correlated distribution. Ideally, from such initializations, the dynamics would escape from the saddle points separating modes quickly and spread its mass between the dominant modes.
@@ 1 Introduction
A ubiquitous and generically hard computational task is to minimize a high-dimensional function $f$ over a discrete space $\{1,...,q\}^{n}$; closely connected is the problem of sampling from the probability distribution with mass proportional to $e^{-\beta f}$ for $\beta$ large. The function $f$ is often viewed as an energy landscape in statistical physics or as a loss function in machine learning, and the $\beta$-large setting is referred to as the low-temperature regime. The difficulty is induced by the possibility of $f$ having several minima with large basins of attraction (or, equivalently, by the multi-modality of the induced distribution); this poses a barrier to traditional optimization/sampling algorithms like gradient descent and Markov chain sampling, at least when initialized from a worst-case state, e.g., in a sub-optimal mode.
@@
Still, a black box approach to these tasks would select the initial state randomly from a product measure or more general high-entropy distribution and would hope that this can circumvent (some of) the bottlenecks in the space. Ideally, a high-entropy initialization spreads its mass across the space in such a way that dynamics are primarily driven by diffusion away from the saddles separating dominant modes, picking the basin to fall into with the correct probability. However, rigorous studies of convergence rates to stationarity from high-entropy initializations are severely lacking and generically difficult.
@@
We focus on the theoretical study of high-entropy initializations in the context of Markov chains for spin system models, like the Ising, Potts, and random-cluster models. In the statistical physics literature, questions of dynamics from high-entropy initializations have a long (empirical) history. A paradigmatic version of this is in the Ising model on the integer lattice graph $\mathbb{Z}^{d}$ at low temperatures, where it is widely expected (though entirely open to prove) that the Glauber dynamics (the natural reversible local Markov chain) initialized from a uniform-at-random assignment of ±$\pm 1$ mixes in polynomial time: see, e.g., the review paper [9] for the rich physics literature about this process. Even in other geometries, for instance on random graphs, bounding the speed of convergence of the low-temperature Ising Glauber dynamics initialized uniformly at random from {±$\{\pm 1\}^{n}$ seems to be mathematically very challenging. (See, e.g., [45, 20, 46, 25] for recent progress on zero-temperature $\beta=\infty$ versions of this problem.) In the special case of the Ising Glauber dynamics on the complete graph, known as the mean-field model, the process reduces perfectly to a 1-dimensional birth-death Markov chain. Here [38, 18] showed the escape from the saddle corresponding to balanced configurations at low temperatures is fast, implying $O(n\log n)$ mixing when initialized from a product of fair coin tosses, despite slow mixing from worst-case.
@@
When one generalizes from the Ising model to the $q$-state Potts model, the above questions become significantly more complicated due to the presence of a higher-dimensional space of spin counts, the possibility of slow mixing at intermediate temperatures, and the lack of symmetry between the modes. The closely related random-cluster model faces similar difficulties, as well as the non-locality of its interactions and update rules. Together, these yield a rich class of models for investigating the above-described expected benefits from high-entropy (e.g., product) initializations to overcome the slow mixing of standard Markov chains.
@@
By contrast, the interest in initializations from product measures (and similar high-entropy distributions), is towards understanding success/failure of black-box approaches that do not require prior knowledge of the target distribution’s modes. In particular, such initializations do not rely on, nor hope for, the convexity of the landscape near the initialization, instead relying on the non-convexity to allow access to multiple modes. We also mention that there are close connections between such high-entropy initializations and simulated annealing and tempering schemes, which we expand on in Section 1.3.
In this paper, we study the convergence of Markov chains in spin systems in the presence of metastability and phase coexistence from high-entropy initializations (predominantly, product measures). By phase coexistence, we mean the presence of multiple modes of roughly equal weights, and by metastability, we mean the presence of multiple modes of vastly different weights; in both situations, Markov chains get trapped for exponential time scales far from stationarity. We study two canonical Markov chains, the Chayes–Machta (CM) dynamics for the mean-field random-cluster model and the Glauber dynamics for the mean-field Potts model. These chains are known to be slow mixing from worst-case initializations for ranges of intermediate, near-critical, and low temperatures. We study their equilibration times in their slow-mixing regimes from product measure initializations and sharply characterize the families of product initializations that lead to optimally fast mixing.
Already on the complete graph, the understanding of mixing from such initializations poses significant challenges, and traditionally, the mean-field setting is a natural starting point for the study of random processes in other families of graphs of significant interest. Our analysis requires a careful understanding of low-dimensional projections of the Markov chains near unstable saddle points of the landscape separating its dominant modes; namely, in regions where the fluctuations of the projected chain compete on the same scale with its drift as well as regions of strict non-convexity. Unlike the case of the mean-field Ising Glauber dynamics where the projection onto the count of plus spins is itself a 1-dimensional birth-and-death Markov chain, in our setting, these projections are not themselves Markov chains. Instead, we approximate the projections of truly high-dimensional Markov chains on a “good” set of configurations by more tractable 1-dimensional chains whose diffusion away from saddle points separating modes we then study. Altogether, this amounts to a significantly more refined analytic control of the Markov chains than previous works on these dynamics. See Section 1.4 for more on the proof ideas.
@@ 1.1 Main results
The (ferromagnetic) Potts model on the $n$-vertex complete graph $G=([n],\binom{n}{2})$ at inverse temperature $\beta>0$ is the probability distribution $\pi_{\beta}$ over spin assignments $\sigma\in\{1,..,q\}^{n}$ to vertices of $G$ such that
$\pi_{\beta}(\sigma)=\frac{1}{Z_{\beta,q}}\cdot\exp\Big{(}-\frac{\beta}{n}\sum_{i\neq j}\mathbf{1}\{\sigma_{i}\neq\sigma_{j}\}\Big{)}\,,$ (1)
where $Z_{\beta}$ is the normalizing constant known as the partition function and $\sigma_{i}$ denotes the spin of vertex $i$.
@@
$\mu_{\beta}(A)=\frac{1}{Z_{p,q}}\cdot\Big{(}\frac{p}{1-p}\Big{)}^{|A|}q^{k([n],A)}\,,$ (2)
where $k([n],A)$ is the number of connected components in the subgraph induced by $A$, and $Z_{p,q}$ is the corresponding partition function. Note that when $q=1$, the mean-field random-cluster model corresponds exactly to the Erdős–Rényi random graph model, but when $q\neq 1$ the cluster weighting can change the phenomenology significantly. When $q\geq 2$ is integer and $p=1-e^{-\beta/n}$, the random-cluster model is closely linked to the Potts model; in particular, if one assigns spins to the components of $A\sim\mu_{\beta}$ independently and uniformly at random among $\{1,...,q\}$, the result is a sample from the Potts distribution at inverse temperature $\beta$. To unify the discussion, via the reparametrization $p=1-e^{-\beta/n}\approx\beta/n$, we henceforth only work with a temperature parameter $\beta$, even when discussing the random-cluster model.
@@
The standard way to quantify the speed of convergence of a Markov chain is the mixing time, i.e., the time it takes to reach total-variation distance $\varepsilon$ to stationarity, either from a prescribed initialization, or from a worst-case initialization. For the worst-case mixing, as soon as $\varepsilon<1/2$, the TV-distance decays exponentially fast, so one usually takes as a convention $\varepsilon=1/4$. In the context of this paper, we are considering mixing times from a specified initial distribution, and often in settings where the worst-case initialization mixing time is exponentially large. This does not satisfy the same boosting property (i.e., exponential decay of TV-distance once it is $<1/2$) as the worst-case mixing time does, so when we refer to bounds on mixing time from an initial distribution, we mean that in that time, we can attain any TV-distance $\varepsilon>0$ independent of $n$. Note that when we discuss mixing from an initial distribution, the law of the process is over both the initialization and the run of the Markov chain.
To describe what is known about the mixing time behavior of the CM and Glauber dynamics, we first recall the static phase diagram of the mean-field Potts and random-cluster models. These have been extensively analyzed and are controlled by three threshold values $\beta_{\mathsf{u}}\leq\beta_{\mathsf{c}}\leq\beta_{\mathsf{s}}$ of $\beta$. The middle one, $\beta_{\mathsf{c}}$, is the critical point marking the distinction between order (a dominant spin class in the Potts setting or, equivalently, a “giant” connected component in the random-cluster) and disorder (balanced spin classes or small connected components) in a typical sample from the distribution. When $q>2$, $\beta_{\mathsf{u}}$ and $\beta_{\mathsf{s}}$ are two additional critical points marking the onset of metastability for the ordered phases and the end of metastability of the disordered phase, respectively (see Figure 2) and [6, 15, 40].
@@
Figure 1: The negative log probability (i.e., free energy) of the largest connected component size in the $n$-vertex mean-field random-cluster model with $q=3$ (for large $n$). The $x$-axes are the fraction of vertices in the largest component, and the $y$-axes represent the logarithm of the total weight of the configurations with such largest component sizes divided by $n$. We emphasize that the landscape is a function of all the component sizes and can exhibit more complexity in the other directions, but our results show this 1-dimensional projection approximately governs the dynamics from product initializations.
@@
We start by stating our main result for the CM dynamics demonstrating the sharp families of product initializations that circumvent the exponential slowdown throughout the metastability regime $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. Let $\bigotimes\text{Ber}(\lambda_{0}/n)$ be the product distribution where each edge is included with probability $\lambda_{0}/n$. (Note that $\bigotimes\text{Ber}(\lambda_{0}/n)$ corresponds to the classical Erdős–Rényi $G(n,\lambda_{0}/n)$ random graph model.)
@@
Recall that the CM dynamics is fast from every initialization when $\beta\notin(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ per [4, 2], so this covers all temperatures where the worst-case mixing time is slow. The constant $\lambda_{*}$ is explicit and is such that the expected size of the giant in an Erdős–Rényi $(n,\lambda_{*}/n)$ is $\theta_{*}n$ where $\theta_{*}$ is the saddle point in the landscape of the random-cluster model projected onto its giant component size, separating the ordered and disordered phases: see Figure 1. The constant $c_{*}$ is defined in terms of left vs. right exit probabilities for a 1-dimensional Gaussian process. This is described detail in a detailed proof overview in Section 1.4.
As hinted by fact that $\lambda_{*}$ places the giant exactly at the saddle point separating the ordered and disordered phases, the thresholds of Theorem 1.1 are sharp. Namely, if $\lambda_{0}$ is not in the delineated regimes, then there is an $\varepsilon>0$ such that the $\varepsilon$-mixing time is exponentially slow: see Section 1.2. This demonstrates that even from the perspective of product measure initializations, it is still important that some parameter of the product measure be tuned according to the model from which one is sampling.
@@
We now turn to our results for the Potts Glauber dynamics that show the sharp families of high-entropy initializations that overcome the exponential bottlenecks in the metastable and phase-coexistence regimes. Let $\hat{\nu}^{\otimes}(m_{0})$ denote the distribution over $\{1,...,q\}^{[n]}$, where first, a dominant spin $s$ is selected uniformly at random from $\{1,\dots,q\}$, and then, independently, each vertex is assigned spin $s$ with probability $m_{0}\geq 1/q$ and each of the remaining $q-1$ spins with probability $\frac{1-m_{0}}{q-1}$.
@@
When $\beta\leq\beta_{\mathsf{u}}$ the mixing time is fast from arbitrary initialization per [16], so the above covers all temperature regimes where the worst-case mixing time is slow. The constant $m_{*}$ is explicit and corresponds to initializing exactly at the saddle point(s) separating the $q$ ordered phases from the disordered phase (see Figure 2). We note that in the low-temperature $\beta>\beta_{\mathsf{s}}$ regime, our result includes the special case $m_{0}=1/q$, which is the fully uniform-at-random assignment of spins to vertices, and requires understanding the delicate competition between multiple simultaneously unstable directions around the saddle. Finally, the ranges of $m_{0}$ in the theorem statement are sharp for fast mixing from initializations of the form $\hat{\nu}^{\otimes}$; the following subsection expands on what we mean by this sharpness.
@@
Figure 2: The negative log probability (i.e., free energy) of the $n$-vertex mean-field Potts model with $q=3$, projected down onto the fractional spin counts; i.e., the $xy$-axes are the fraction of vertices assigned spin $1$ and $2$ respectively (which also determines the fraction of vertices with spin $3$), and the $z$-axis is the negative logarithm of the configuration weights divided by $n$. The orange-marked spots are the saddle points near which Glauber dynamics may be initialized in Theorem 1.2.
1.2 Sharpness of the initialization parameter regimes
The parameter values $\lambda_{0}$ and $m_{0}$ in Theorems 1.11.2 are such that in Figures 12, the initializations can be microscopically near the saddle, so long as they are on the correct side of it. We also obtain slow mixing even from the families of high-entropy initializations considered, when the initialization parameters are not in the windows in Theorems 1.11.2. Namely, for every $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$,
  • Theorem 3.24 establishes that the CM dynamics are exponentially slow to attain $o(1)$ TV-distance to stationarity when initialized from $\bigotimes\text{Ber}(\lambda_{0}/n)$ for $\lambda_{0}$ outside the range specified in Theorem 1.1;
  • Theorem 4.22 establishes that Potts dynamics are exponentially slow to attain $o(1)$ TV-distance to stationarity when initialized from $\hat{\nu}^{\otimes}(m_{0})$ for $m_{0}$ outside the range specified in Theorem 1.2.
These results are a consequence of our proof strategy, which shows that whichever side of the saddles at $\lambda_{*}$ or $m_{*}$ the initailization starts on, it quasi-equilibrates to the corresponding phase, and then takes exponential time to leave that phase. E.g., if $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$ and $\lambda_{0}<\lambda_{*}+O(n^{-1/2})$ then the CM dynamics has an $\Omega(1)$ chance of quasi-equilibrating to the disordered phase. This implies slow mixing when combined with the fact that for $\beta\in(\beta_{\mathsf{c}},\beta_{\mathsf{s}})$, initialized from the disordered phase, mixing is exponentially slow.

### v2

@@ Abstract
A common obstruction to efficient sampling from high-dimensional distributions with Markov chains is the multimodality of the target distribution because they may get trapped far from stationarity. Still, one hopes that this is only a barrier to the mixing of Markov chains from worst-case initializations and can be overcome by choosing high-entropy initializations, e.g., a product or weakly correlated distribution. Ideally, from such initializations, the dynamics would escape from the saddle points separating modes quickly and spread its mass between the dominant modes with the correct probabilities.
@@ 1 Introduction
A ubiquitous and generically hard computational task is to minimize a high-dimensional function $f$ over a discrete space $\{1,...,q\}^{n}$; closely connected is the problem of sampling from the probability distribution with mass proportional to $e^{-\beta f}$ for $\beta$ large. The function $f$ is often viewed as an energy landscape in statistical physics or as a loss function in machine learning, and the $\beta$-large setting is referred to as the low-temperature regime. The difficulty is induced by the possibility of $f$ having several minima with large basins of attraction (or, equivalently, by the multimodality of the induced distribution); this poses a barrier to traditional optimization/sampling algorithms like gradient descent and Markov chain sampling, at least when initialized from a worst-case state, e.g., in a sub-optimal mode.
@@
Still, a black box approach to these tasks would select the initial state randomly from a product measure or more general high-entropy distribution and would hope that this can circumvent (some of) the bottlenecks in the space. When we say high-entropy initialization, we mean distributions that are well-spread over the probability space, in contrast to, e.g., worst-case initializations, or initializations in some extremal energy state. Ideally, a high-entropy initialization distributes its mass across the space in such a way that dynamics are primarily driven by diffusion away from the saddles separating dominant modes, picking the basins to fall into with the correct probability. (Here and throughout we are using the terminology “saddle point" informally, by analogy to critical points of landscapes in continuous spaces.) However, rigorous studies of convergence rates to stationarity from high-entropy initializations are severely lacking and generically difficult.
@@
We focus on the theoretical study of high-entropy initializations in the context of Markov chains for spin system models, such as the Ising, Potts, and random-cluster models. In the statistical physics literature, questions of dynamics from high-entropy initializations have a long (empirical) history. A paradigmatic version of this is in the Ising model on the integer lattice graph $\mathbb{Z}^{d}$ at low temperatures, where it is widely expected (though entirely open to prove) that the Glauber dynamics (the natural reversible local Markov chain) initialized from a uniform-at-random assignment of $\pm 1$ mixes in polynomial time: see, e.g., the review paper [9] for the rich physics literature about this process. Indeed, this question requires understanding the motion of interfaces separating regions where plus and minus respectively dominate, and is likely even harder than the notorious problem of showing a polynomial bound for the worst-case initialization mixing time of Ising Glauber dynamics in a box with plus boundary conditions [45, 46, 42].
Even in other geometries, for instance on trees and random graphs, bounding the speed of convergence of the low-temperature Ising Glauber dynamics initialized uniformly at random from $\{\pm 1\}^{n}$ seems to be mathematically very challenging. (See, e.g., [37, 50, 21, 51, 26] for recent progress on zero-temperature $\beta=\infty$ versions of this problem, and [11] for analysis on the tree initialized from a biased product measure.) In the special case of the Ising Glauber dynamics on the complete graph, known as the mean-field model, the process reduces perfectly to a 1-dimensional birth-death Markov chain. Here [40, 19] showed the escape from the saddle corresponding to balanced configurations at low temperatures is fast, implying $O(n\log n)$ mixing when initialized from a product of fair coin tosses, despite slow mixing from worst-case.
@@
When one generalizes from the Ising model to the $q$-state Potts model, the above questions become significantly more complicated due to the presence of a higher-dimensional space of spin counts, the possibility of slow mixing at intermediate temperatures, and the lack of symmetry between the modes. The closely related random-cluster model faces similar difficulties, as well as the non-locality of its interactions and update rules. Together, these yield a rich class of models for investigating the above-described expected benefits and possible limitations for high-entropy (e.g., product) initializations to overcome the slow mixing of standard Markov chains.
@@
By contrast, the interest in initializations from product measures (and similar high-entropy distributions), is towards understanding success/failure of black-box approaches that do not require prior knowledge of the target distribution’s modes. In particular, such initializations do not rely on, nor hope for, local convexity of the landscape near the initialization, instead relying on the non-convexity to allow access to multiple modes, making the analysis significantly more challenging. We also note the close connections between high-entropy initializations and simulated annealing schemes, for which we obtain new results: see Section 1.2.
In this paper, we study the convergence of Markov chains in spin systems in the presence of metastability and phase coexistence from high-entropy initializations (predominantly, product measures). By phase coexistence, we mean the presence of multiple modes of roughly equal weights, and by metastability, we mean the presence of multiple modes of vastly different weights; in both situations, Markov chains get trapped for exponential time scales far from stationarity. We study two canonical Markov chains, the Chayes–Machta (CM) dynamics for the mean-field random-cluster model and the Glauber dynamics for the mean-field Potts model; our results also extend to the well-studied Swendsen-Wang (SW) dynamics. These chains are known to be slow mixing from worst-case initializations for ranges of intermediate, near-critical, and low temperatures. We study their equilibration times in their slow-mixing regimes from product measure initializations and sharply characterize the families of product initializations that lead to optimally fast mixing.
Already on the complete graph, the understanding of mixing from such initializations poses significant challenges. Our analysis requires a careful understanding of low-dimensional projections of the Markov chains near unstable saddle points of the landscape separating its dominant modes; namely, in regions where the fluctuations of the projected chain compete on the same scale with its drift as well as regions of strict non-convexity. Unlike the case of the mean-field Ising Glauber dynamics where the projection onto the count of plus spins is itself a 1-dimensional birth-and-death Markov chain, in our setting, these projections are not themselves Markov chains. Instead, we approximate the projections of truly high-dimensional Markov chains on a “good” set of configurations by more tractable 1-dimensional chains whose diffusion away from saddle points separating modes we then study. Altogether, this amounts to a significantly more refined analytic control of the Markov chains than previous works on these dynamics. See Section 1.3 for more on the proof ideas.
@@ 1.1 Main results
The (ferromagnetic) Potts model on the $n$-vertex complete graph $G=([n],\binom{[n]}{2})$ at inverse temperature $\beta>0$ is the probability distribution $\pi_{\beta}$ over spin assignments $\sigma\in\{1,\dots,q\}^{n}$ to vertices of $G$ such that
$\pi_{\beta}(\sigma)=\frac{1}{Z_{\beta,q}}\cdot\exp\Big{(}\frac{\beta}{n}\sum_{i\neq j}\mathbf{1}\{\sigma_{i}=\sigma_{j}\}\Big{)}\,$, (1)
where $Z_{\beta,q}$ is the normalizing constant known as the partition function and $\sigma_{i}\in\{1,\dots,q\}$ denotes the spin of vertex $i$.
@@
$\mu_{p}(A)=\frac{1}{Z_{p,q}}\cdot\Big{(}\frac{p}{1-p}\Big{)}^{|A|}q^{k([n],A)}\,$, (2)
where $k([n],A)$ is the number of connected components in the subgraph induced by $A$, and $Z_{p,q}$ is the corresponding partition function. Note that when $q=1$, the mean-field random-cluster model corresponds exactly to the Erdős–Rényi random graph model, but when $q\neq 1$ the cluster weighting can change the phenomenology significantly. When $q\geq 2$ is integer and $p=1-e^{-\beta/n}$, the random-cluster model is closely linked to the Potts model; in particular, if one assigns spins to the components of $A\sim\mu_{\beta}$ independently and uniformly at random among $\{1,...,q\}$, the result is a sample from the Potts distribution at inverse temperature $\beta$. To unify the discussion, via the reparametrization $p=1-e^{-\beta/n}\approx\beta/n$, we henceforth only work with a temperature parameter $\beta$, even when discussing the random-cluster model, and we write $\mu_{\beta}$ for $\mu_{p}$.
@@
The standard way to quantify the speed of convergence of a Markov chain is the mixing time, i.e., the time it takes to reach total-variation distance $\varepsilon$ to stationarity, either from a prescribed initialization, or from a worst-case initialization. For the worst-case mixing, as soon as $\varepsilon<1/2$, the TV-distance decays exponentially fast, so one usually takes as a convention $\varepsilon=1/4$. In the context of this paper, we are considering mixing times from a specified initial distribution, and often in settings where the worst-case initialization mixing time is exponentially large. This does not satisfy the same boosting property (i.e., exponential decay of TV-distance once it is <1/2<1/2) as the worst-case mixing time does, so when we refer to bounds on mixing time from an initial distribution, we mean that in that time, we can attain any TV-distance $\varepsilon>0$ independent of $n$. Note that when we discuss mixing from an initial distribution, the law of the process is over both the initialization and the run of the Markov chain.
To describe what is known about the mixing time behavior of the CM and Glauber dynamics, we first recall the static phase diagram of the mean-field Potts and random-cluster models. These have been extensively analyzed and are controlled by three threshold values $\beta_{\mathsf{u}}\leq\beta_{\mathsf{c}}\leq\beta_{\mathsf{s}}$ of $\beta$. The middle one, $\beta_{\mathsf{c}}$, is the critical point marking the distinction between order (a dominant spin class in the Potts setting or, equivalently, a “giant” connected component in the random-cluster) and disorder (balanced spin classes or small connected components) in a typical sample from the distribution. When $q>2$, $\beta_{\mathsf{u}}$ and $\beta_{\mathsf{s}}$ are two additional critical points marking the onset of metastability for the ordered phases and the end of metastability of the disordered phase, respectively (see Figure 1, Figure 2 and [6, 16, 43].)
@@
Figure 1: The negative log probability (i.e., restricted free energy) of the largest connected component’s size in the $n$-vertex mean-field random-cluster model with $q=3$ (for large $n$). The $x$-axes are the fraction of vertices in the largest component, and the $y$-axes represent the logarithm of the total weight of the configurations with such largest component sizes divided by $n$. We emphasize that the landscape is a function of all the component sizes and can exhibit more complexity in the other directions, but our results show this 1-dimensional projection approximately governs the dynamics from product initializations. The orange-marked dots are the saddle points with respect to which the fast initializations for Chayes–Machta dynamics are characterized in Theorem 1.1
@@
We start by stating our main results for the CM dynamics demonstrating the sharp families of product initializations that circumvent the exponential slowdown throughout the metastability regime $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. Let $\bigotimes\text{Ber}(\lambda_{0}/n)$ be the product distribution where each edge is included with probability $\lambda_{0}/n$. (Note that $\bigotimes\text{Ber}(\lambda_{0}/n)$ corresponds to the classical Erdős–Rényi $G(n,\lambda_{0}/n)$ random graph model.)
@@
Conversely, if $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ and the CM dynamics is initialized from $\bigotimes\text{Ber}(\lambda_{0}/n)$ for $\lambda_{0}$ outside the corresponding regime of 13 (meaning, e.g., for item 1., that $\lambda_{0}\geq\lambda_{*}(\beta,q)-O(n^{-1/2})$), then it takes $\exp(\Omega(n))$ time to attain any $o(1)$ total-variation distance. [^1] By this we mean that there exists an $\varepsilon_{0}$ and a $C_{0}$ such that attaining any $\varepsilon<\varepsilon_{0}$ total-variation distance takes time $\exp(C_{0}n)$.
[…]
The most delicate part of the analysis is the $\beta=\beta_{c}$ case, where the coexistence of ordered and disordered phases necessitates pinning down the relative exit probabilities to the right and left of the saddle from initializations whose giant component size is within order of the standard deviation from (in particular, microscopically near) the saddle point. This is the source of the constant $c_{*}$, which is defined in terms of left vs. right exit probabilities for an explicit approximating 1-dimensional Gaussian process. This is also in contrast to sampling schemes that use mixtures of extremal initializations [27, 22, 28] where one starts a priori by matching the initialization weights with the weights of the different phases at stationarity. This is fleshed out in a detailed proof overview in Section 1.3.
@@
We now turn to our results for the Potts Glauber dynamics that show the sharp families of high-entropy initializations that overcome the exponential bottlenecks in the metastable and phase-coexistence regimes. Let $\hat{\nu}^{\otimes}(m_{0})$ denote the distribution over $\{1,...,q\}^{[n]}$, where first, a dominant spin $s$ is selected uniformly at random from $\{1,\dots,q\}$, and then, independently, each vertex is assigned spin $s$ with probability $m_{0}\geq 1/q$ and each of the remaining q1q-1 spins with probability $\frac{1-m_{0}}{q-1}$.
@@
Conversely, if $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ and the Potts Glauber dynamics is initialized from $\hat{\nu}^{\otimes}(m_{0})$ for $m_{0}$ outside the corresponding regime of 13 (meaning, e.g., for item 1., that $m_{0}\geq m_{*}(\beta,q)-O(n^{-1/2})$), then it takes $\exp(\Omega(n))$ time to attain any $o(1)$ total-variation distance.
When $\beta\leq\beta_{\mathsf{u}}$ the mixing time is fast from arbitrary initializations per [17], so the above covers all temperature regimes where the worst-case mixing time is slow. The constant $m_{*}$ is explicit and corresponds to initializing exactly at the saddle point(s) separating the $q$ ordered phases from the disordered phase (see Figure 2). This again is the source of the sharpness of the classification. We note that in the low-temperature $\beta>\beta_{\mathsf{s}}$ regime, our result includes the most delicate and important case $m_{0}=1/q$, which is the fully uniform-at-random assignment of spins to vertices; this requires understanding the subtle competition between multiple simultaneously unstable directions around the saddle.
Remark 2.
The critical initialization parameters $c_{*},\hat{c}_{*}$ are defined by the points at which the approximating 1D Markov chains have certain (explicit in terms of $q$) left and right exit probabilities. If one is interested in computationally finding $c_{*},\hat{c}_{*}$ this can be done up to the requisite precision $o(1/\sqrt{n})$ as follows. Since the exit probabilities are smooth and monotone in the initialization (see proofs of Lemma 3.22 and Lemma 4.15), perform binary search on the potential value for $c_{*},\hat{c}_{*}$, with $\tilde{O}(n)$ many runs of the 1D Markov chain to estimate the exit probabilities at each search point. This leads to a complexity of $\tilde{O}(n)$ for estimating the parameters.
Remark 3.
A lower bound of $\Omega(n\log n)$ on the mixing time of the Glauber Potts dynamics will generally hold, except for the potentially delicate case, corresponding to the one special value of $m_{0}$ where the initialization puts the fractional size of the largest color class exactly at the stable fixed point $m_{r}$. For other values of $m_{0}$, an $\Omega(n\log n)$ lower bound follows from the concentration estimates and bounds on the derivative of the drift function as it approaches the fixed point found in Section 4. We predict that the lower bound exactly at the $m_{r}$ initialization is instead $\Omega(n)$, but it is somewhat subtle. (Similarly, an $\Omega(\log n)$ lower bound holds for the CM dynamics except potentially at the one special value of $\lambda_{0}$ where it places the fractional giant size exactly at $\theta_{r}$.)
@@
Figure 2: The negative log probability (i.e., free energy) of the $n$-vertex mean-field Potts model with $q=3$, projected down onto the fractional spin counts; i.e., the xyxy-axes are the fraction of vertices assigned spin 11 and 22 respectively (which also determines the fraction of vertices with spin 33), and the $z$-axis is the negative logarithm of the configuration weights divided by $n$. The orange-marked dots are the saddle points with respect to which the fast initializations for Glauber dynamics are characterized in Theorem 1.2.